OpenStudy (anonymous):

There are 6 pounds of potatoes. Tammy needs eight-ninths of a pound of potatoes to make one serving of a dish. How many servings can she make?

OpenStudy (stephstar101):

6/(8/9)=x, solve for x

OpenStudy (stephstar101):

I'll tell you if you're right or not :D

OpenStudy (anonymous):

\[\frac{8}{9}*x=6\]

OpenStudy (anonymous):

six and three-fourths?

OpenStudy (stephstar101):

Yup!! Good job!

OpenStudy (anonymous):

\[\text{Floor}\left[\frac{6}{\frac{8}{9}}\right]=6 \]

OpenStudy (anonymous):

\[\frac{6}{\frac{8}{9}}=\frac{27}{4}=6.75 \]6 whole servings